# Transfer matrix approach to quantum systems subject to certain Lindblad evolution

### v1

@@ V Conclusion
In conclusion, we demonstrated the use of transfer matrix approach to solve for the two point correlators of an XX spin chain with homogeneous dephasing. We showed that this approach leads to a simple solution of all the elements of the correlation matrix in terms of products of $2\times 2$ transfer matrices. From the solution, the thermodynamic limit follows directly due to the explicit presence of $L$ dependence in the products. Our main result in Eq. (40) provides a simple expression for the evolution of the magnetization density at any time, $t$ and from this expression the asymptotic behaviors at short and long times of the magnetization density become readily apparent. We also looked at the asymptotic behaviors of the off-diagonal elements and found that at long times they can be expressed as higher order spatial derivatives of the diagonal components. The latter gives that the long time behavior of the $l^{th}$ off-diagonal elements is given by $1/t^{\lceil l/2\rceil+1}$.
While we only looked at a rather simple model with local dephasing, this approach can in principle be utilized to solve for the two point correlators of other quadratic Hamiltonians subject to Hermitian Lindblad operators that give a closed set of linear equations for the correlators, respectively. Clearly, for complicated Lindblad operators it may not be possible to get as simple expressions as Eq. (40). However, we can always obtain the transfer matrix equation of the type in Eq. (28) which, as we demonstrated, can be used numerically to determine the evolution of density. The complexity of the transfer matrix method with transfer matrices of dimension $r$ is of the order of $L^{2}r^{6}$ as opposed to exact diagonalization or iterations of Eq. (10) which scales as $L^{6}$. Therefore, one is able to access large system sizes as long as the transfer matrices are of small dimension. Their dimensions depend on the range of the couplings present in the system, and therefore will be small as long as the system has just nearest neighbor or next-to-nearest neighbor coupling. For example, the non-local dephasing models Žnidarič (2024); Wang et al. (2024) with three site dephasing operators has a $4\times 4$ transfer matrix.
@@ Appendix A Even-Odd structure of the correlation matrix
Here we show using induction that if the initial correlation matrix, $C_{x,y}(0)$, is such that its even off-diagonal elements are real and odd off-diagonal elements are imaginary then it remains so for all the times. Assuming that it is true at time $t_{n}$, then at the next step
@@ Appendix B Derivation of the Transfer Matrix
$T^{L-1}=\frac{1}{\sin(z)}\begin{pmatrix}\sin(zL)&-\sin(z(L-1))\\ \sin(z(L-1))&-\sin(z(L-2))\end{pmatrix},$ (60)
where $z=\arccos[-iu].$
@@ Appendix C Laplace Inverse
where $\eta$ is taken to be bigger than real parts of all the singularities of $(\sqrt{s^{2}+\omega^{2}(q)}-4\gamma)^{-1}$. This integral can be computed using a Bromwich contour where the analytic structure of the integrand in the complex $s$ plane becomes important. As stated in the main text, the integrand has a branch cut which we choose to run between the two branch points $b_{\pm}=\pm i\omega(q)$ (See Fig (1)). The integrand also has poles located at $s_{\pm}=\pm\sqrt{4\gamma^{2}-\omega^{2}(q)}$. The poles lie on the real axis for $4\gamma>\omega(q)$ and on the imaginary axis for $4\gamma<\omega(q)$. Since $\omega(q)=8\sin(q/2)$ is $q$ dependent, the two cases of real and imaginary poles needs to be considered separately.
@@
where $I_{r}(t,q)$, $I_{i}(t,q)$ are given by equations Eq. (63) and Eq. (64), respectively.
One can quickly see that in the long time limit the exponential damping, $e^{-4\gamma t}$, kills off all the terms except the long wavelength contribution from the pole on the real axis i.e. the first term in $I_{r}(t,q)$ at small $q$. Therefore, dropping all other terms and expanding the first term in $I_{r}(t,q)$ around $q=0$, we recover the diffusion kernal. To get the short time behavior requires some algebra. To that end, rewriting the integrals in $I_{r}(t,q)$ and $I_{i}(t,q)$ in terms of the variable $s=\omega\cos\theta$, the resulting expressions for $C_{x,x}(t)$ become the same as in Ref. Eisler, 2011 upto a re-parametrization. One can then do similar analysis as in Ref. Eisler, 2011 to recover the result in Eq. (42).

### v2

@@ V Conclusion
In conclusion, we demonstrated the use of transfer matrix approach to solve for the two point correlators of an XX spin chain with homogeneous dephasing. We showed that this approach leads to a simple solution for all the elements of the correlation matrix in terms of products of $2\times 2$ transfer matrices. From the solution, the thermodynamic limit follows directly due to the explicit presence of $L$ dependence in the products. Our main result in Eq. (42) provides a simple expression for the evolution of the magnetization density at any time, $t$ and from this expression the asymptotic behaviors at short and long times of the magnetization density become readily apparent. We also looked at the asymptotic behaviors of the off-diagonal elements and found that at long times they can be expressed as higher order spatial derivatives of the diagonal components. The latter gives that the long time behavior of the $l^{th}$ off-diagonal elements is given by $1/t^{\lceil l/2\rceil+1/2}$.
While we only looked at a rather simple model with local dephasing, this approach can in principle be utilized to solve for the two point correlators of other quadratic Hamiltonians subject to Hermitian Lindblad operators that give a closed set of linear equations for the correlators. Clearly, for complicated Lindblad operators it may not be possible to get as simple expressions as Eq. (42). However, we can always obtain the transfer matrix equation of the type in Eq. (30) which, as we demonstrated, can be used numerically to determine the evolution of the density. The complexity of the transfer matrix method with transfer matrices of dimension $r$ is of the order of $L^{2}r^{6}$ as opposed to exact diagonalization or iterations of Eq. (10) which scales as $L^{6}$. Therefore, one is able to access large system sizes as long as the transfer matrices are of small dimension. The dimensions of the transfer matrix depend on the range of the couplings present in the system, and therefore will be small as long as the system has just nearest neighbor or next-to-nearest neighbor coupling. For example, the non-local dephasing models Žnidarič (2024); Wang et al. (2024) with three site dephasing operators has a $4\times 4$ transfer matrix.
@@ Appendix A Even-Odd structure of the correlation matrix
Here we show that if the initial correlation matrix, $C_{x,y}(0)$, is such that its even off-diagonal elements are real and odd off-diagonal elements are imaginary then it remains so for all the times. Assuming that it is true at time $t_{n}$, then at the next step
@@ Appendix B Derivation of the Transfer Matrix
$T^{L-1}=\frac{1}{\sin(\alpha)}\begin{pmatrix}\sin(\alpha L)&-\sin(\alpha(L-1))\\ \sin(\alpha(L-1))&-\sin(\alpha(L-2))\end{pmatrix},$ (63)
where $\alpha=\arccos[-iu].$
@@ Appendix C Laplace Inverse
where $\eta$ is taken to be bigger than real parts of all the singularities of $(\sqrt{s^{2}+\omega^{2}(q)}-4\gamma)^{-1}$. This integral can be computed using a Bromwich contour where the analytic structure of the integrand in the complex $s$ plane becomes important. As stated in the main text, the integrand has a branch cut which we choose to run between the two branch points $b_{\pm}=\pm i\omega(q)$ (See Fig (1)). The integrand also has poles located at $s_{\pm}=\pm\sqrt{4\gamma^{2}-\omega^{2}(q)}$. The poles lie on the real axis for $4\gamma>\omega(q)$ and on the imaginary axis for $4\gamma<\omega(q)$. Since $\omega(q)=8J\sin(q/2)$ is $q$ dependent, the two cases of real and imaginary poles needs to be considered separately.
@@
where $I_{r}(t,q)$, $I_{i}(t,q)$ are given by Eq. (66) and Eq. (67), respectively.
One can quickly see that in the long time limit the exponential damping, $e^{-4\gamma t}$, kills off all the terms except the long wavelength contribution from the pole on the real axis i.e. the first term in $I_{r}(t,q)$ at small $q$. Therefore, dropping all other terms and expanding the first term in $I_{r}(t,q)$ around $q=0$, we recover the diffusion kernel. To get the short time behavior requires some algebra. To that end, rewriting the integrals in $I_{r}(t,q)$ and $I_{i}(t,q)$ in terms of the variable $s=\omega\cos\theta$, the resulting expressions for $C_{x,x}(t)$ become the same as in Ref. Eisler, 2011 upto a re-parametrization. One can then do similar analysis as in Ref. Eisler, 2011 to recover the result in Eq. (45).